OpenStudy (anonymous):

find the derivative of the function. f(x)=(u^2-9)/(u^2+9) from [2x,3x]

OpenStudy (zehanz):

f is a function of x, but the formula has u in it...

OpenStudy (anonymous):

oh my bad i left out the du at the end. f(x)=(u^2-9)/(u^2+9)du

OpenStudy (zehanz):

Shouldn't there also be an integral sign in there somewhere?

OpenStudy (abb0t):

Do you mean find the anti-derivative?

OpenStudy (anonymous):

yeah theres an integral sign sorry

OpenStudy (zehanz):

So\[f(x)=\int\limits_{2x}^{3x}\frac{ u^2-9 }{ u^2+9 }du\]And you need f'(x), right?

OpenStudy (anonymous):

correct

OpenStudy (zehanz):

Using the Fundamental Theorem of Calculus: \[f(x)=G(3x)-G(2x)\]where G is the antiderivative of the integrand. The integrand being\[g(u)=\frac{ u^2-9 }{ u^2+9 }\] Now all you have to do is:\[f'(x)=\left( G(3x)-G(2x) \right)'=...\]

OpenStudy (zehanz):

This turns into:\[f'(x)=(G(3x))'-(G(2x))'=3\cdot g(3x)-2 \cdot g(2x)=...\]So just plug in u=3x and u=2x in g(u) to get f'(x).